A group of people want to share equally the cost of a $200,000 airplane. They want to find 5 more people to join their club so that the cost per person decreases by $2000. How many people are currently in the club?
Answer by checkley71(8403)

200,000/X=200,000/(X+5)+2000
200,000/X-200,000/(X+5)=2,000 COMBINE THE X FRACTIONS USING A COMMON DENOMINATOR.
[200,000(X+5)-200,000X)/X(X+5)=2,000
(200,000X+1,000,000-200,000X)/(X^2+5X)=2,000 NOW CROSS MULTIPLY
2,000(X^2+5X)=1,000,000
2,000(X^2+5X)-1,000,000=0
2,000(X^2+5X-500)=0
2,000(X+25)(X-20)=0
X-20=0
X=20 IS THE NUMBER OF ORIGINAL CLUB MEMBERS.
PROOF:
200,000/20=200,000/25+2,000
10,000=8,000+2,000
10,000=10,000